OpenStudy (anonymous):

What is the value of 12C4 ? A) 11,880 B) 792 C) 495 D) 48

OpenStudy (raden):

use the formula : nCr = n!/r!(n-r)! in this case n=12 and r=4

OpenStudy (anonymous):

so would it be 12*11*10*9?

OpenStudy (raden):

then divide by 4!

OpenStudy (raden):

4! = 4 * 3 * 2 * 1 = 24

OpenStudy (anonymous):

oh ok i think i got it! kinda haha ,do you have time for another one?

OpenStudy (raden):

i'll try if i can :)

OpenStudy (anonymous):

B

OpenStudy (raden):

you are wrong, BigNasty have you done before ?

OpenStudy (anonymous):

Kathi can choose 1 snack from a box containing 6 different granola bars, 4 different kinds of fruit, and 2 different kinds of chips. How many ways can Kathi choose a snack? a)100 b) 26 c) 20 d) 12

OpenStudy (anonymous):

It's an attempt

OpenStudy (raden):

addition rule : 6+4+2=12

OpenStudy (raden):

because just choose 1 snack

OpenStudy (anonymous):

yep i got that one corrrecwhat about this one? What is the value of (8 - 4)! ? a)6 24 5040 40,296

OpenStudy (raden):

8-4=4 so, 4! = 4 * 3 * 2 * 1 = ...

OpenStudy (anonymous):

24